OpenStudy (anonymous):

The length of a rectangle is 3cm more than twice its width. Find the largest possible width if the perimeter is at most 66cm. use an inequality to solve.

OpenStudy (anonymous):

So where are you stuck? How did you set up the equation for perimeter?

OpenStudy (anonymous):

i just don't get it

OpenStudy (anonymous):

can you walk me through it? I will give you best response

OpenStudy (anonymous):

thanks molly :) you will be getting best response

OpenStudy (anonymous):

Well, here's what we know. Perimeter = 2 * length + 2* height. And L = 2*W + 3 And of course, P < = 66.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so what next?

OpenStudy (anonymous):

So, just plugging the 2W+3 equivalent into the formula for perimeter we get P = 2*(2*W + 3) + 2*W must be less than or equal to 66.

OpenStudy (anonymous):

so let y represent the length and x represent width. y=2x+3 Perimeter=2x+2y Substitute in 2x+3 for y (P=2x+2(2x+3)) So now, P=6x+6 Plug in 66 for p and make it an inequality 66\[66\ge6x+6\] then you would sovlve and get x=10

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

sorry \[x \le10\]